### v2

# Bridging the Linguistic Divide: A Survey on Leveraging Large Language Models for Machine Translation

Abstract
The advent of Large Language Models (LLMs) has significantly reshaped the landscape of machine translation (MT), particularly for low-resource languages and domains that lack sufficient parallel corpora, linguistic tools, and computational infrastructure. This survey presents a comprehensive overview of recent progress in leveraging LLMs for MT. We analyze techniques such as few-shot prompting, cross-lingual transfer, and parameter-efficient fine-tuning (e.g., LoRA, adapters) that enable effective adaptation to under-resourced settings. The paper also explores synthetic data generation strategies using LLMs, including back-translation and lexical augmentation. Additionally, we compare LLM-based translation with traditional encoder-decoder models across diverse language pairs, highlighting the strengths and limitations of each. We discuss persistent challenges—such as hallucinations, evaluation inconsistencies, and inherited biases—while also evaluating emerging LLM-driven metrics for translation quality. This survey offers practical insights and outlines future directions for building robust, inclusive, and scalable MT systems in the era of large-scale generative models.
This survey provides a comprehensive exploration of how Large Language Models (LLMs) are transforming machine translation (MT). By systematically analyzing techniques such as few-shot prompting, cross-lingual transfer, and parameter-efficient fine-tuning, the study highlights practical methods to enhance translation quality despite data scarcity. It also identifies critical challenges, including hallucinations, biases, and evaluation bottlenecks. The insights from this work can guide researchers and practitioners in developing more robust and scalable MT systems, driving inclusivity for underrepresented languages and improving cross-lingual communication.
Large Language Models, Machine Translation, Neural Machine Translation
1 Introduction
Machine Translation (MT) has significantly evolved from early rule-based systems to advanced neural models. Initially, MT relied heavily on manually crafted linguistic rules and dictionaries, requiring extensive human intervention for each language pair and struggling with linguistic nuances, idiomatic expressions, and contextual understanding [1]. The transition to Statistical Machine Translation (SMT) in the late 1980s and early 1990s marked a major advancement, as SMT utilized probabilistic models trained on large bilingual corpora, automating translation processes [2, 3]. However, SMT faced challenges such as limited fluency, difficulties with less common language pairs, and issues handling long-range dependencies. The introduction of neural networks marked a turning point in MT research. Recurrent Neural Networks (RNNs) and later, sequence-to-sequence models with Long Short-Term Memory (LSTM) [4] units improved translation fluency by learning contextual dependencies [5, 6]. However, these architectures faced limitations in handling long sequences due to vanishing gradients and computational constraints. The emergence of attention mechanisms [7, 8] addressed these issues by allowing models to focus selectively on relevant parts of the input sequence, enhancing both accuracy and coherence. This led to the development of the Transformer model [9], which became the foundation for modern neural MT systems.
Despite advancements in MT, translating low-resource languages remains a significant challenge. Most high-performing MT models rely on extensive parallel corpora, which are scarce for many languages. Researchers have explored various strategies, including transfer learning [10, 11, 12, 13, 14], multilingual training [15], data augmentation [16, 17], etc., to improve translation quality for low-resource languages. These approaches have shown promise but still face difficulties in capturing linguistic diversity, domain adaptation, and ensuring consistent translation quality.
Challenges in MT extend beyond data scarcity. Linguistic ambiguity, polysemy [18], domain-specific terminology, and cultural nuances [19, 20] continue to pose obstacles to accurate translation. Additionally, hallucination, where the model generates plausible yet incorrect [21, 22], and robustness against adversarial attacks [23, 24, 25] remain active areas of research.
Large Language Models (LLMs) offer new opportunities to address these challenges. Their ability to learn from vast amounts of multilingual and monolingual text allows them to generate translations with improved fluency and contextual awareness. Few-shot and zero-shot learning capabilities enable LLMs to handle low-resource languages with minimal supervision. Moreover, LLMs facilitate domain adaptation and personalized translation by leveraging in-context learning. However, they also introduce concerns such as increased computational costs, potential biases, and the necessity for efficient deployment strategies.
As MT continues to advance, the integration of LLMs is expected to redefine the field, offering new possibilities while necessitating careful consideration of ethical and computational trade-offs. Future research must focus on improving efficiency, addressing bias, and ensuring reliable translations across diverse languages and domains.
2 Structure of the Survey
This survey provides a comprehensive analysis of the role of Large Language Models (LLMs) in Machine Translation (MT), particularly focusing on low-resource languages. The paper is structured as follows:
Section III: Synthetic Data Generation with LLMs – This section explores how LLMs can be leveraged to generate synthetic bilingual data, a critical technique for improving MT performance in data-scarce scenarios. It reviews various strategies, including back-translation, forward translation, and domain adaptation techniques.
Section IV: Translation Methodologies – Here, we categorize MT approaches using LLMs into two primary paradigms: (1) Prompting-Based Techniques, which leverage pre-trained LLMs without modifying their parameters, and (2) Fine-Tuning Approaches, which adapt LLMs to specific translation tasks by modifying their weights. We provide an in-depth analysis of methods such as zero-shot prompting, in-context learning, supervised fine-tuning, and parameter-efficient fine-tuning techniques like LoRA.
Section V: LLMs vs Encoder-Decoder across Languages – This section evaluates how LLMs perform across different languages, with a special focus on low-resource language pairs.
Section VI: LLM based Evaluation – The final technical section discusses various evaluation metrics and methodologies used to assess LLM-driven MT.
This study offers a comprehensive analysis of how large language models (LLMs) are transforming machine translation (MT), particularly in low-resource settings where parallel data is limited. It highlights the emerging challenges and opportunities, providing a foundation for future research in this rapidly evolving domain. To ensure methodological consistency, the study refrains from direct cross-paper comparisons, as existing works vary significantly in terms of datasets, language pairs, and evaluation metrics.
For clarity and focus, the analysis is restricted to text-to-text, sentence-level translation, excluding document-level, dialogue-based, and multimodal approaches (e.g., image or video-assisted translation, sign language translation). While certain studies employ hybrid techniques—such as reinforcement learning (RL) combined with quantized low-rank adaptation (QLoRA)—they are categorized under the primary methodology most representative of their core contribution.
Given the vast and diverse landscape of MT research, this work specifically concentrates on sentence-level translation, aiming to provide a precise and coherent overview of the latest advancements driven by LLMs.
3 Synthetic Data Generation with LLMs
Domain adaptation in Machine Translation (MT) has been explored by leveraging large pre-trained language models (LMs) for synthetic data generation [26]. This approach addresses two scenarios: one with limited bilingual in-domain data and another with no bilingual in-domain data. In the former case, target-side synthetic text is generated using GPT-J (English) and mGPT (Arabic), followed by back-translation to create parallel data. In the latter case, source text is forward-translated using an MT model before applying the same augmentation technique. A mixed fine-tuning strategy with in-domain data oversampling is employed to ensure balanced learning, resulting in significant BLEU [27, 28] score improvements of approximately 5-6 for Arabic-English and 2-3 for English-Arabic. The improvement in quality has been also confirmed by manual evaluations.
Another synthetic data generation technique, LexMatcher, focuses on addressing missing senses of polysemous words [29]. By using bilingual dictionaries, this method retrieves data from parallel corpora to include existing senses. Subsequently, data augmentation is performed with an LLM to generate new sentence pairs for uncovered word senses, enhancing translation disambiguation and sense coverage.
The use of word and phrase alignment for synthetic data generation has also been studied [30]. However, findings indicate that synthetic data produced by LLMs does not offer significant improvements over rule-based and neural-based approaches [31].
4 Translation Methodologies
Machine Translation using LLMs can be broadly categorized into two primary methodological approaches: (1) prompting-based techniques and (2) fine-tuning-based techniques. The former leverages pre-trained models with minimal adaptation, whereas the latter involves modifying model parameters to enhance performance on translation tasks. The methods can be combined - prompting based techniques can be used after fine-tuning. In this section, we discuss these methodologies in detail.
4.1 Prompting-Based Techniques
Prompting leverages pre-trained LLMs by conditioning them on task-specific instructions or examples without modifying their underlying parameters. This approach is particularly useful when computational resources for fine-tuning are limited.
4.1.1 Zero-Shot Prompting
In zero-shot prompting, the LLM is provided with an instruction to translate a sentence without any examples. The model relies solely on its pre-trained knowledge to generate translations.
Different approaches in zero-shot prompting vary by prompt format and the use of a pivot language for low-resource languages. It has been observed that zero-shot prompting with ChatGPT lags behind MT systems by Google MT [32], Tencent [33], and DeepL by around 5.0 BLEU points [34]. Pivot prompting has been explored to translate between distant languages, where the LLM first translates the sentence to English and then into the target language. This strategy, which uses a resource-rich language (English) as a pivot, improves translation quality between DeZh and RoZh [34]. Human evaluation on the generated translations reveals that GPT-4 makes fewer mistakes compared to ChatGPT and Google MT. Interestingly, although GPT-4 scores lower in BLEU compared to Google MT, human annotators rank its translations more favorably in blind evaluations [34].
To further enhance zero-shot translation, additional information such as domain-specific context and keywords has been incorporated into the prompts. Including domain information leads to improved translation quality compared to vanilla zero-shot prompting, while providing around 10 keywords related to the domain achieves even better COMET scores [35, 36]. Similarly, adding POS tags in the prompt has been shown to generate slightly better quality translations [37].
Variations in Prompt Template
Prompts that specify only the source and target language have been found to achieve the best overall results, using a simple format [38] such as:
[src_lang]: input
[tgt_lang]:
Similarly, the authors in [34] tested three prompt variations and the following format was found to yield the best performance:
Please provide the [tgt_lang]
translation for these sentences:
Using pre-trained models for translation without instruction tuning poses challenges, as the model often generates responses that are diverse and difficult to parse. For instance, it sometimes continues generating tokens even after completing the translation [39, 40]. A common solution is to discard tokens after a newline; however, models frequently produce outputs in patterns that do not match known criteria, making this strategy unreliable. These issues are generally mitigated when models undergo fine-tuning [39, 40].
4.1.2 In-context Learning
In-Context Learning (ICL) enhances machine translation by providing a model with a small set of input-output examples before generating a translation, helping it infer correct patterns. This method significantly improves translation accuracy, especially for domain-specific or idiomatic expressions, by guiding the model toward appropriate mappings. The effectiveness of the approach depends on multiple parameters, including the number of in-context examples, selection of the examples, and their ordering.
However, the use of in-context examples does not always result in better output quality. It has been observed that in some cases, few-shot prompting can lead to poorer performance compared to one-shot prompting [39]. When models are fine-tuned without few-shot examples, they exhibit a higher likelihood of hallucination when prompted with few-shot examples during inference. Fine-tuning with few-shot examples has been shown to mitigate this issue, reducing the chances of hallucination and enhancing output quality.
Number of In-Context Examples: One critical decision when designing prompts for large language models (LLMs) is determining the optimal number of in-context examples. Increasing the number of examples can enhance performance, but it also comes with higher computational costs and diminishing returns. Using too few examples may bias the model towards the style of the provided instances, while incorporating too many may not lead to proportional improvements in output quality.
Several studies have explored this trade-off. Research has shown consistent improvements in COMET scores when increasing from one to five in-context examples [41]. Similarly, a significant rise in BLEU scores has been reported when increasing from a single example to five, although results for intermediate number of examples were not provided [42].
[…]
These findings collectively suggest that while increasing the number of in-context examples enhances performance, there exists a threshold beyond which further improvements are marginal.
Selection of In-Context Examples: Selecting effective in-context examples is crucial for optimizing translation performance in LLMs. While strategic selection of examples can enhance results, studies suggest that random selection may sometimes be just as effective, provided the examples are of high quality [43]. The overall quality of examples, regardless of whether they are randomly or carefully chosen, plays a significant role in performance.
The impact of noisy or irrelevant examples has also been investigated. Research indicates that even a single poor-quality example can severely degrade translation quality, particularly in 1-shot scenarios [46]. Increasing the number of examples can mitigate the negative effects of one bad sample, but optimal performance is achieved when selecting high-quality, relevant exemplars. Notably, a single optimized example has been found to outperform multiple random ones. Additionally, semantic similarity between the input and in-context examples has been positively correlated with translation quality [44].
One method for optimizing example selection is CTQScorer, which refines prompt selection by incorporating multiple linguistic and structural features [47]. This approach starts with BM25 retrieval to generate a candidate pool and then ranks examples based on factors such as semantic similarity, translation quality, sentence length, and perplexity. By leveraging a contextual translation quality score, this method has demonstrated a 2.5-point improvement in COMET score over random selection.
Further advancing example selection strategies, research has highlighted the importance of coherence in in-context learning [48]. Rather than relying solely on similarity-based retrieval, coherence is defined as the semantic and stylistic consistency between prompt examples and the test sentence. Studies show that ensuring coherence—aligning prompts in meaning, style, and contextual flow—yields better translation performance than selecting examples purely based on word or embedding similarity. This approach shifts the focus from isolated, highly similar examples to logically consistent prompt setups, redefining how in-context examples should be chosen for machine translation.
An alternative approach involves using submodular functions to optimize example selection [49]. This technique ensures that adding a new example improves coverage while avoiding excessive redundancy. By balancing content overlap, target alignment, and diversity, submodular-selected prompts have consistently outperformed random and purely similarity-based selections across multiple datasets, leading to superior few-shot translation results.
In low-resource scenarios, selecting source-similar examples has been found to yield better BLEU and COMET scores than using arbitrary examples [50]. When data is scarce, closely matched examples can provide valuable lexical and structural guidance that the model might not otherwise recall. Additionally, the makeup of the candidate pool influences selection effectiveness—when limited to high-quality reference translations, the performance gap between random and similarity-based selection decreases.
The authors in [42] used translation memories (TMs) to enhance large language model (LLM) translation performance. The approach retrieves the top-k most similar sentence pairs from a TM database and incorporates them into prompts using structured templates. Two main template styles—instruction-style and code-style—are used to provide translation hints. Experiments demonstrate that TM-based prompting significantly improves LLM translation quality, often surpassing strong NMT baselines.
The importance of correctly pairing in-context examples has also been emphasized [51]. Mismatched examples, such as incorrect translations, can lead to complete translation failures. Interestingly, cross-lingual demonstrations have been shown to be surprisingly effective in low-resource translation scenarios. Using examples from a different but better-known language pair has, in some cases, outperformed using same-language examples. This suggests that LLMs may leverage general translation patterns across linguistically related languages, improving translation quality even when direct examples in the target language are unavailable.
Collectively, these findings underscore the importance of intelligent example selection—considering factors such as relevance, quality, coherence, and diversity—to maximize few-shot translation performance.
Example Ordering:
The ordering of examples in the prompt dataset plays a important role in determining the quality of translations. This section explores how dataset ordering can be optimized to enhance translation performance. Using a BERT-based re-ranker model to rank the examples has been shown to improve BLEU [27], ROUGE [52], CHRF [53], and COMET [54] scores [45]. Similarly, examining the effect of example order in few-shot prompts reveals that task-level prompts are less sensitive to ordering compared to randomly sampled examples, resulting in lower standard deviations and better translation quality. Testing all permutations of four randomly sampled and top task-level examples demonstrates that the ordering of in-context examples significantly impacts MT performance, as indicated by notable BLEU variations across three runs with random examples [46].
Experiments with two types of ordering, least similar to most similar (followed by the input sentence) and most similar to least similar (followed by the input sentence), show no significant improvements of either method over the other. However, the former yields slightly higher average BLEU and COMET scores [50].
4.2 Fine-Tuning Approaches
Fine-tuning involves modifying the parameters of an LLM to improve its translation capabilities. This method is computationally expensive but results in a model that is better adapted to specific translation tasks. Common fine-tuning techniques include:
4.2.1 Supervised Fine-Tuning (SFT)
Supervised fine-tuning (SFT) involves training an LLM on a large parallel corpus by minimizing the cross-entropy loss between predicted translations and ground truth references. This method significantly enhances the model’s translation accuracy, especially for specialized domains. It has been shown that SFT enables LLMs to achieve strong multilingual translation performance with as few as 32 examples, where training on one language direction generalizes to others [55]. However, using English as the target language can lead to misinterpretations, and the presence of target-side noise tends to affect well-represented languages more adversely. Ensuring successful SFT alignment requires the model to maintain a surface-level focus, preventing it from acquiring unwanted biases during the fine-tuning process.
4.2.2 Low-Rank Adaptation (LoRA)
LoRA [56] is a parameter-efficient fine-tuning technique that reduces memory overhead by injecting small trainable matrices into frozen pre-trained models. This approach enables effective adaptation to new translation domains without modifying the entire model. A comprehensive empirical study on parameter-efficient fine-tuning (PEFT) methods for machine translation, conducted across various parameter budgets and language pairs, observed that using adapters achieves performance on par with full fine-tuning while training only 10% of the parameters [57]. However, the effectiveness of PEFT methods varies significantly based on language pair distance and dataset size, with noticeable struggles for distant language pairs. We refer the readers to [58] and [59] for a detailed analysis for PEFT methods.
The QLoRA method [60] has been applied to fine-tune Mistral-7B [61], achieving better results than GPT-3.5-turbo when combined with one-shot prompting [62]. In a comparison between LoRA and supervised fine-tuning (SFT), both methods were found to deliver similar performance, although LoRA requires training 50 times fewer parameters [39]. Furthermore, it was demonstrated that LoRA can surpass the few-shot performance of pretrained models with as few as 2,000 training examples.
An evaluation of 15 publicly available language models on machine translation tasks showed that QLoRA fine-tuning outperforms both few-shot learning and models trained from scratch, achieving a significant BLEU score improvement while training on only 0.77% of the model parameters [63]. Adapter-based fine-tuning with LoRA was further shown to match the performance of traditional fine-tuning, while reducing the number of training parameters by a factor of 50 [39].
4.2.3 ALMA
A two-stage fine-tuning approach has been introduced to enhance machine translation capabilities in large language models (LLMs) [64]. Since many LLMs are predominantly trained on English data, the first stage, known as ALMA, fine-tunes the model on non-English monolingual datasets. This step strengthens the model’s proficiency in various languages, enabling a more robust understanding of linguistic nuances across different tongues.
In the subsequent stage, ALMA fine-tunes the model using a relatively small set of high-quality parallel data. This phase focuses on translation generation, aligning the model’s outputs with accurate translations and improving its performance on translation tasks. The two-stage fine-tuning strategy has demonstrated significant improvements, achieving an average increase of over 12 BLEU and 12 COMET scores compared to the model’s zero-shot performance across 10 translation directions from the WMT’21 and WMT’22 test datasets.
4.2.4 Reinforcement Learning based Finetuning
The standard SFT methodology, which focuses on reducing the gap between predicted outputs and gold-standard references, inherently limits model performance to the quality of the training data. This presents a notable challenge, as even human-authored data—typically regarded as high-quality—can contain imperfections. To address this limitation, the authors in [65] introduced Contrastive Preference Optimization (CPO) for machine translation. CPO differentiates itself by training models to distinguish between “adequate” and “near-perfect” translations. This is achieved through a contrastive learning strategy that presents the model with hard negative examples—translations that are high-quality but not flawless. By learning to reject these near-perfect yet imperfect translations, the model is encouraged to produce superior outputs. When applied to ALMA models with only 22K parallel sentences and fine-tuning only 0.1% of the parameters, CPO yielded significant improvements. The authors in [66] uses Reinforcement Learning with Human Feedback (RLHF) to improve machine translation by aligning model outputs with human preferences. Instead of relying on expensive human-annotated preference data, the authors use high-quality human translations from books as a proxy for preferred outputs. The process involves fine-tuning a base translation model on parallel corpora, training a reward model to distinguish between machine and human translations, and then refining the translation model through RL using Proximal Policy Optimization (PPO) [67]. This approach enhances translation quality and shows cross-lingual transferability, meaning the improvements generalize to languages not directly optimized with RLHF. The results demonstrate that RLHF can effectively capture human translation preferences in a cost-efficient manner.
5 LLMs vs Encoder-Decoder across Languages
This section explores how large language models (LLMs) perform across languages and how they compare to dedicated machine translation (MT) systems. An early study found that ChatGPT lags behind traditional MT systems such as Google Translate, DeepL, and Tencent, but observed that GPT-4 performs competitively with these systems in zero-shot settings [34]. This indicates that some LLMs achieve performance on par with traditional MT systems for high-resource languages.
When translating from Spanish to 11 indigenous South American languages, fine-tuned M2M100 was shown to consistently outperform GPT-4 [68]. The performance gap further increases when M2M100 is combined with techniques like kNN-MT [69]. The contrast between high-resource and low-resource languages is also highlighted by a multitask evaluation, where ChatGPT’s performance, although slightly lower, remains close to fine-tuned state-of-the-art models [70]. For instance, ChatGPT achieves 58.64 chrF++ compared to 63.5 for fine-tuned models in XXEn translation, and 51.12 chrF++ versus 54.4 for EnXX [71]. However, for low-resource languages, the difference becomes much more pronounced, with ChatGPT scoring 27.79 chrF++ compared to 54.9 for fine-tuned models in XXEn, and 21.57 versus 41.9 chrF++ for EnXX [53].
6 LLM based Evaluation
The performance of large language models (LLMs) on the MQM-2020 dataset has been evaluated by comparing their Spearman correlations with traditional machine translation (MT) metrics, such as ROUGE, BERTScore, and BARTScore [72]. The findings indicate that, when provided with appropriate instructions and demonstrations, LLMs can outperform these conventional metrics. However, the comparison is restricted to high-resource language pairs (English and Chinese), and more advanced metrics, such as COMET, which exhibit higher correlation with human judgments, were not included.
For MT evaluation, GEMBA-style prompting [73] has been employed, showing that GPT-4 achieves the highest system-level pairwise accuracy and the highest Kendall’s tau [74] for English-to-German (EnDe) and English-to-Russian (EnRu) translations [75]. In contrast, COMET demonstrates superior performance for Chinese-to-English (ZhEn), achieving the highest Kendall’s tau correlation.
Building on these approaches, Error Analysis Prompting has been introduced to combine error analysis with chain-of-thought reasoning for MT evaluation using LLMs [76]. Experiments reveal that, while LLMs achieve higher system-level accuracy, they fall short of traditional metrics, including embedding-based approaches such as BLEURT, COMET, and MetricX, at the segment level.
The distinction between system-level and segment-level evaluation is crucial in MT assessment. System-level evaluation refers to the overall performance of an MT system, typically measured by aggregating scores across multiple sentences or documents. In contrast, segment-level evaluation assesses the quality of individual translated sentences, offering a more fine-grained analysis. While LLMs demonstrate strong system-level accuracy, their lower segment-level accuracy suggests potential inconsistencies when evaluating individual translations, highlighting a key limitation compared to traditional metrics.
7 Summary and Recommendation
This survey has provided a comprehensive analysis of how Large Language Models (LLMs) are reshaping the landscape of machine translation (MT). The review highlighted the effectiveness of LLMs in generating synthetic data, leveraging few-shot and in-context learning, and fine-tuning through parameter-efficient techniques such as LoRA and QLoRA. While LLMs show competitive performance in high-resource language pairs, they still lag behind traditional MT systems for low-resource languages due to data scarcity and linguistic diversity challenges. Additionally, the evaluation of LLM-based MT demonstrates that although LLMs can achieve competitive system-level accuracy, they face limitations in segment-level consistency compared to traditional metrics like COMET.
Enhanced Adaptation for Low-Resource Languages: Despite advancements, LLMs struggle with low-resource languages. Future research should focus on improved transfer learning strategies, cross-lingual pretraining, and multilingual fine-tuning to enhance performance in low-data settings. Leveraging smaller, domain-specific datasets and semi-supervised learning could boost generalization in low-resource scenarios.
Reinforcement Learning for Quality Control: The application of reinforcement learning techniques (e.g., RLHF or Contrastive Preference Optimization) can refine LLMs’ output by aligning them with human preferences. This would reduce hallucination issues and improve translation reliability, particularly in domain-specific and low-resource contexts.
Better Evaluation and Interpretability: Future work should explore more robust and interpretable evaluation metrics. Incorporating explainability techniques, such as attention visualization or error analysis prompting, could help in understanding LLM translation errors and guiding model improvements.
Acknowledgment
Baban Gain and Dibyanayan Bandyopadhyay gratefully acknowledge the Prime Minister’s Research Fellowship (PMRF) program for providing financial support and enabling this research.