### v3

# Bridging the Linguistic Divide: A Survey on Leveraging Large Language Models for Machine Translation

Abstract
The advent of Large Language Models (LLMs) has significantly reshaped the landscape of machine translation (MT), particularly for low-resource languages and domains that lack sufficient parallel corpora, linguistic tools, and computational infrastructure. This survey presents a comprehensive overview of recent progress in leveraging LLMs for MT. We analyze techniques such as few-shot prompting, cross-lingual transfer, and parameter-efficient fine-tuning (e.g., LoRA, adapters) that enable effective adaptation to under-resourced settings. The paper also explores synthetic data generation strategies using LLMs, including back-translation and lexical augmentation. Additionally, we compare LLM-based translation with traditional encoder-decoder models across diverse language pairs, highlighting the strengths and limitations of each. We discuss persistent challenges—such as hallucinations, evaluation inconsistencies, and inherited biases—while also evaluating emerging LLM-driven metrics for translation quality. This survey offers practical insights and outlines future directions for building robust, inclusive, and scalable MT systems in the era of large-scale generative models.
This survey provides a comprehensive exploration of how Large Language Models (LLMs) are transforming machine translation (MT). By systematically analyzing techniques such as few-shot prompting, cross-lingual transfer, and parameter-efficient fine-tuning, the study highlights practical methods to enhance translation quality despite data scarcity. It also identifies critical challenges, including hallucinations, biases, and evaluation bottlenecks. The insights from this work can guide researchers and practitioners in developing more robust and scalable MT systems, driving inclusivity for underrepresented languages and improving cross-lingual communication.
Large Language Models, Machine Translation, Neural Machine Translation
1 Introduction
Machine Translation (MT) has significantly evolved from early rule-based systems to advanced neural models. Initially, MT relied heavily on manually crafted linguistic rules and dictionaries, requiring extensive human intervention for each language pair and struggling with linguistic nuances, idiomatic expressions, and contextual understanding [1]. The transition to Statistical MT (SMT) in the late 1980s and early 1990s marked a major advancement, as SMT utilized probabilistic models trained on large bilingual corpora, automating translation processes [2, 3]. However, SMT faced challenges such as limited fluency, difficulties with less common language pairs, and issues handling long-range dependencies. The introduction of neural networks marked a turning point in MT research. Recurrent Neural Networks (RNNs) and later, sequence-to-sequence models with Long Short-Term Memory (LSTM) [4] units improved translation fluency by learning contextual dependencies [5, 6]. However, these architectures faced limitations in handling long sequences due to vanishing gradients and computational constraints. The emergence of attention mechanisms [7, 8] addressed these issues by allowing models to focus selectively on relevant parts of the input sequence, enhancing both accuracy and coherence. This led to the development of the Transformer model [9], which became the foundation for modern neural MT systems. MT is now widely integrated into global communication—from business collaborations to cross-cultural teamwork enabling multilingual interactions in real time or asynchronously. A study [10] show that MT-mediated exchanges significantly enhance communication clarity, discussion depth, and decision-making quality in multilingual team settings.
Despite advancements in MT, translating low-resource languages remains a significant challenge. Most high-performing MT models rely on extensive parallel corpora, which are scarce for many languages. Researchers have explored various strategies, including transfer learning [11, 12, 13, 14, 15], multilingual training [16], data augmentation [17, 18], etc., to improve translation quality for low-resource languages. These approaches have shown promise but still face difficulties in capturing linguistic diversity, domain adaptation, and ensuring consistent translation quality.
Challenges in MT extend beyond data scarcity. Linguistic ambiguity, polysemy [19], domain-specific terminology, and cultural nuances [20, 21] continue to pose obstacles to accurate translation. Additionally, hallucination, where the model generates plausible yet incorrect [22, 23], and robustness against adversarial attacks [24, 25, 26] remain active areas of research.
LLMs offer new opportunities to address these challenges. Their ability to learn from vast amounts of multilingual and monolingual text allows them to generate translations with improved fluency and contextual awareness. Few-shot and zero-shot learning capabilities enable LLMs to handle low-resource languages with minimal supervision. Moreover, LLMs facilitate domain adaptation and personalized translation by leveraging in-context learning. However, they also introduce concerns such as increased computational costs, potential biases, and the necessity for efficient deployment strategies.
As MT continues to evolve, the integration of LLMs is poised to fundamentally transform the field—introducing new capabilities while also raising important ethical and computational considerations. Although prior surveys have reviewed various aspects of MT [27, 28, 29], to the best of our knowledge, none have systematically captured the substantial advancements enabled by LLMs in recent years. Given the diverse array of techniques available—ranging from prompting strategies to parameter-efficient fine-tuning—navigating this emerging landscape can be challenging for newcomers and practitioners alike. To address this gap, this paper presents a comprehensive analysis of the principal methodologies for leveraging LLMs in MT, offering detailed comparisons and insights to guide future research and application.
2 Structure of the Survey
This survey provides a comprehensive analysis of the role of LLM in MT, particularly focusing on low-resource languages. The paper is structured as follows:
Section III: Synthetic Data Generation with LLMs – This section explores how LLMs can be leveraged to generate synthetic bilingual data, a critical technique for improving MT performance in data-scarce scenarios. It reviews various strategies, including back-translation, forward translation, and domain adaptation techniques.
Section IV: Translation Methodologies – Here, we categorize MT approaches using LLMs into two primary paradigms: (1) Prompting-Based Techniques, which leverage pre-trained LLMs without modifying their parameters, and (2) Fine-Tuning Approaches, which adapt LLMs to specific translation tasks by modifying their weights. We provide an in-depth analysis of methods such as zero-shot prompting, in-context learning, supervised fine-tuning, and parameter-efficient fine-tuning techniques like LoRA.
Section V: LLMs vs Encoder-Decoder across Languages – This section evaluates how LLMs perform across different languages, with a special focus on low-resource language pairs.
Section VI: LLM based Evaluation – The final technical section discusses various evaluation metrics and methodologies used to assess LLM-driven MT.
This study offers a comprehensive analysis of how LLMs are being used for MT. It highlights the emerging challenges and opportunities, providing a foundation for future research in this rapidly evolving domain. To ensure methodological consistency, the study refrains from direct cross-paper comparisons, as existing works vary significantly in terms of datasets, language pairs, and evaluation metrics.
For clarity and focus, the analysis is restricted to text-to-text, sentence-level translation, excluding document-level[30, 31], dialogue-based, and multimodal approaches [32] (e.g., image or video-assisted translation, sign language translation [33]). While certain studies employ hybrid techniques—such as reinforcement learning (RL) combined with quantized low-rank adaptation (QLoRA)—they are categorized under the primary methodology most representative of their core contribution.
Given the vast and diverse landscape of MT research, this work specifically concentrates on sentence-level translation, aiming to provide a precise and coherent overview of the latest advancements driven by LLMs.
3 Synthetic Data Generation with LLMs
Domain adaptation in MT has been explored by leveraging pre-trained LLMs for synthetic data generation [34]. This approach addresses two scenarios: one with limited bilingual in-domain data and another with no bilingual in-domain data. In the former case, target-side synthetic text is generated using GPT-J [35] (English) and mGPT [36] (Arabic), followed by back-translation [37] to create parallel data. In the latter case, source text is forward-translated using an MT model before applying the same augmentation technique. A mixed fine-tuning strategy with in-domain data oversampling is employed to ensure balanced learning, resulting in significant BLEU [38, 39] score improvements of approximately 5-6 for Arabic-English and 2-3 for English-Arabic. The improvement in quality has been also confirmed by manual evaluations.
Another synthetic data generation technique, LexMatcher, focuses on addressing missing senses of polysemous words [40]. By using bilingual dictionaries, this method retrieves data from parallel corpora to include existing senses. Subsequently, data augmentation is performed with an LLM to generate new sentence pairs for uncovered word senses, enhancing translation disambiguation and sense coverage.
The use of word and phrase alignment for synthetic data generation has also been studied [41]. However, findings indicate that synthetic data produced by LLMs does not offer significant improvements over rule-based and neural-based approaches [42], although the study focused on only one language. On a broader study, covering more languages, the authors in [43] generate a document-level synthetic corpus by forward-translating English Europarl data into seven diverse low-resource languages, later expanding it to 147 additional language pairs via pivoting. They evaluate the quality of this data using automatic metrics and human annotation, confirming its general reliability. Crucially, they show that MT models trained solely on this synthetic data can rival or outperform baselines trained on real data, and that fine-tuning strong pretrained systems like NLLB [44] and LLaMA-3B [45] on this data yields significant performance improvements. Furthermore, their experiments highlight that even noisy synthetic data can be highly effective when clean alternatives are scarce.
4 Translation Methodologies
MT using LLMs can be broadly categorized into two primary methodological approaches: (1) prompting-based techniques and (2) fine-tuning-based techniques. The former leverages pre-trained models with minimal adaptation, whereas the latter involves modifying model parameters to enhance performance on translation tasks. The methods can be combined - prompting based techniques can be used after fine-tuning. In this section, we discuss these methodologies in detail.
4.1 Prompting-Based Techniques
Prompting leverages pre-trained LLMs by conditioning them on task-specific instructions or examples without modifying their underlying parameters. This approach is particularly useful when computational resources for fine-tuning are limited.
4.1.1 Zero-Shot Prompting
In zero-shot prompting, the LLM is provided with an instruction to translate a sentence without any examples. The model relies solely on its pre-trained knowledge to generate translations.
Different approaches in zero-shot prompting vary by prompt format and the use of a pivot language for low-resource languages. It has been observed that zero-shot prompting with ChatGPT lags behind MT systems by Google MT [46], Tencent [47], and DeepL by around 5.0 BLEU points [48]. Pivot prompting has been explored to translate between distant languages, where the LLM first translates the sentence to English and then into the target language. This strategy, which uses a resource-rich language (English) as a pivot, improves translation quality between DeZh and RoZh [48]. Human evaluation on the generated translations reveals that GPT-4 [49] makes fewer mistakes compared to ChatGPT and Google MT. Interestingly, although GPT-4 scores lower in BLEU compared to Google MT, human annotators rank its translations more favorably in blind evaluations [48].
To further enhance zero-shot translation, additional information such as domain-specific context and keywords has been incorporated into the prompts. Including domain information leads to improved translation quality compared to vanilla zero-shot prompting, while providing around 10 keywords related to the domain achieves even better COMET scores [50, 51]. Similarly, adding POS tags in the prompt has been shown to generate slightly better quality translations [52].
Variations in Prompt Template
Prompts that specify only the source and target language have been found to achieve the best overall results, using a simple format [53] such as:
[src_lang]: input
[tgt_lang]:
Similarly, the authors in [48] tested three prompt variations and the following format was found to yield the best performance:
Please provide the [tgt_lang]
translation for these sentences:
Using pre-trained models for translation without instruction tuning poses challenges, as the model often generates responses that are diverse and difficult to parse. For instance, it sometimes continues generating tokens even after completing the translation [54, 55]. A common solution is to discard tokens after a newline; however, models frequently produce outputs in patterns that do not match known criteria, making this strategy unreliable. These issues are generally mitigated when models undergo fine-tuning [54, 55].
4.1.2 In-context Learning
In-Context Learning (ICL) enhances MT by providing a model with a small set of input-output examples before generating a translation, helping it infer correct patterns. This method significantly improves translation accuracy, especially for domain-specific or idiomatic expressions, by guiding the model toward appropriate mappings. The effectiveness of the approach depends on multiple parameters, including the number of in-context examples, selection of the examples, and their ordering.
However, the use of in-context examples does not always result in better output quality. It has been observed that in some cases, few-shot prompting can lead to poorer performance compared to one-shot prompting [54]. When models are fine-tuned without few-shot examples, they exhibit a higher likelihood of hallucination when prompted with few-shot examples during inference. Fine-tuning with few-shot examples has been shown to mitigate this issue, reducing the chances of hallucination and enhancing output quality.
Number of In-Context Examples: One critical decision when designing prompts for LLMs is determining the optimal number of in-context examples. Increasing the number of examples can enhance performance, but it also comes with higher computational costs and diminishing returns. Using too few examples may bias the model towards the style of the provided instances, while incorporating too many may not lead to proportional improvements in output quality.
Several studies have explored this trade-off. Research has shown consistent improvements in COMET scores when increasing from one to five in-context examples [56]. Similarly, a significant rise in BLEU scores has been reported when increasing from a single example to five, although results for intermediate number of examples were not provided [57].
[…]
These findings collectively suggest that while increasing the number of in-context examples enhances performance, there exists a threshold beyond which further improvements are marginal.
Selection of In-Context Examples: Selecting effective in-context examples is crucial for optimizing translation performance in LLMs. While strategic selection of examples can enhance results, studies suggest that random selection may sometimes be just as effective, provided the examples are of high quality [58]. The overall quality of examples, regardless of whether they are randomly or carefully chosen, plays a significant role in performance.
The impact of noisy or irrelevant examples has also been investigated. Research indicates that even a single poor-quality example can severely degrade translation quality, particularly in 1-shot scenarios [62]. Increasing the number of examples can mitigate the negative effects of one bad sample, but optimal performance is achieved when selecting high-quality, relevant exemplars. Notably, a single optimized example has been found to outperform multiple random ones. Additionally, semantic similarity between the input and in-context examples has been positively correlated with translation quality [59].
One method for optimizing example selection is CTQScorer, which refines prompt selection by incorporating multiple linguistic and structural features [63]. This approach starts with BM25 retrieval to generate a candidate pool and then ranks examples based on factors such as semantic similarity, translation quality, sentence length, and perplexity. By leveraging a contextual translation quality score, this method has demonstrated a 2.5-point improvement in COMET score over random selection.
Further advancing example selection strategies, research has highlighted the importance of coherence in in-context learning [64]. Rather than relying solely on similarity-based retrieval, coherence is defined as the semantic and stylistic consistency between prompt examples and the test sentence. Studies show that ensuring coherence—aligning prompts in meaning, style, and contextual flow—yields better translation performance than selecting examples purely based on word or embedding similarity. This approach shifts the focus from isolated, highly similar examples to logically consistent prompt setups, redefining how in-context examples should be chosen for MT.
An alternative approach involves using submodular functions to optimize example selection [65]. This technique ensures that adding a new example improves coverage while avoiding excessive redundancy. By balancing content overlap, target alignment, and diversity, submodular-selected prompts have consistently outperformed random and purely similarity-based selections across multiple datasets, leading to superior few-shot translation results.
In low-resource scenarios, selecting source-similar examples has been found to yield better BLEU and COMET scores than using arbitrary examples [66]. When data is scarce, closely matched examples can provide valuable lexical and structural guidance that the model might not otherwise recall. Additionally, the makeup of the candidate pool influences selection effectiveness—when limited to high-quality reference translations, the performance gap between random and similarity-based selection decreases.
The authors in [57] used translation memories (TMs) to enhance LLMs translation performance. The approach retrieves the top-k most similar sentence pairs from a TM database and incorporates them into prompts using structured templates. Two main template styles—instruction-style and code-style—are used to provide translation hints. Experiments demonstrate that TM-based prompting significantly improves LLM translation quality, often surpassing strong NMT baselines.
The importance of correctly pairing in-context examples has also been emphasized [67]. Mismatched examples, such as incorrect translations, can lead to complete translation failures. Interestingly, cross-lingual demonstrations have been shown to be surprisingly effective in low-resource translation scenarios. Using examples from a different but better-known language pair has, in some cases, outperformed using same-language examples. This suggests that LLMs may leverage general translation patterns across linguistically related languages, improving translation quality even when direct examples in the target language are unavailable.
Collectively, these findings underscore the importance of intelligent example selection—considering factors such as relevance, quality, coherence, and diversity—to maximize few-shot translation performance.
Example Ordering:
The ordering of examples in the prompt dataset plays a important role in determining the quality of translations. This section explores how dataset ordering can be optimized to enhance translation performance. Using a BERT-based re-ranker model to rank the examples has been shown to improve BLEU [38], ROUGE [68], CHRF [69], and COMET [70] scores [61]. Similarly, examining the effect of example order in few-shot prompts reveals that task-level prompts are less sensitive to ordering compared to randomly sampled examples, resulting in lower standard deviations and better translation quality. Testing all permutations of four randomly sampled and top task-level examples demonstrates that the ordering of in-context examples significantly impacts MT performance, as indicated by notable BLEU variations across three runs with random examples [62].
Experiments with two types of ordering, least similar to most similar (followed by the input sentence) and most similar to least similar (followed by the input sentence), show no significant improvements of either method over the other. However, the former yields slightly higher average BLEU and COMET scores [66].
4.2 Advanced Prompting
From a Chain-of-Thought (CoT) perspective, one of the early studies investigated whether step-by-step reasoning prompts, successful in many NLP tasks, can enhance ChatGPTś performance in MT [71]. They apply both zero-shot and 1-shot CoT prompting, where ChatGPT is instructed to translate source sentences word-by-word before generating the full translation. However, unlike in reasoning tasks, CoT induces a literal, phrase-by-phrase translation behavior that hampers fluency and semantic coherence, leading to a significant drop in COMET scores (e.g., -8.8 for ENZH). This degradation highlights a misalignment between CoT prompting and the holistic nature of translation, suggesting that while CoT can benefit structured reasoning, it is ill-suited for MT directly.
[…]
The MAPS (Multi-Aspect Prompting and Selection) [74] framework introduces a prompting strategy that mimics human translation behavior by guiding LLMs through a structured, three-step process. First, the LLM is prompted to extract translation-relevant knowledge from the source sentence in three forms—keywords, the underlying topic, and relevant example translations—using manually designed few-shot prompts. Next, each type of extracted knowledge is separately reintegrated into new prompts to guide the LLM in generating multiple translation candidates. Finally, a selection mechanism, either via a reference-free quality estimation model (COMET-QE) or prompting the LLM itself (LLM-SCQ), is used to choose the best candidate. This multi-faceted prompting strategy improves translation quality across various language pairs and also helps to reduce common errors such as hallucination, ambiguity, and mistranslation without relying on external data or curated in-context examples.
Another method which use Recognizing that translations generated by LLMs often contain subtle errors despite strong benchmark performance, the authors in [75] propose TEaR - a plug-and-play self-refinement framework for MT that improves translation quality through internal feedback loops. TEaR operates in three stages: Translate, where an LLM produces an initial translation; Estimate, where the same LLM simulates human-like error assessment using MQM guidelines; and Refine, where the LLM revises the translation based on its self-generated evaluation. This self-contained design allows TEaR to bypass reliance on external models or human feedback. The Estimate module identifies major and minor translation errors—such as mistranslation, awkward style, and grammar issues—and only proceeds to refinement if necessary, making the approach more selective and stable than baselines like CoT. Experiments on WMT22 and WMT23 datasets across eight translation directions show that TEaR consistently improves translation quality and reducing MQM error counts more effectively than other self-refinement or external feedback systems. Moreover, iterative application of TEaR further enhances translation quality in successive rounds, demonstrating its capacity for continuous improvement.
While prior works attempt to mimic human cognitive behaviors in prompting—such as self-reflection, iterative refinement, or chain-of-thought reasoning—to improve model performance, these methods often fall short when it comes to generating truly divergent or corrective reasoning paths. In particular, self-reflection-based approaches can suffer from what the authors term the Degeneration-of-Thought (DoT) problem: once an LLM gains confidence in its initial (possibly incorrect) answer, it becomes resistant to revising or challenging that stance in subsequent iterations. To address this, the paper proposes the Multi-Agent Debate (MAD) framework [76], which draws on the human problem-solving strategy of adversarial discussion. MAD simulates a structured debate between multiple LLM agents—typically an affirmative and a negative debater—who challenge each other in a tit for tat style, with a judge model overseeing the debate and determining the final answer.
This debate setup introduces external feedback and perspective shifts that self-reflection lacks, promoting more diverse and accurate reasoning. The MAD framework is evaluated on two challenging tasks—Commonsense MT and Counter-Intuitive Arithmetic Reasoning—where it significantly outperforms baseline methods including self-reflection and even matches or surpasses stronger models like GPT-4 when used with GPT-3.5 as the base. Key to MAD’s effectiveness are its adaptive break strategy (which ends debates early when a satisfactory solution is found), balanced argumentative intensity, and consistent use of the same LLM for all roles to mitigate judge bias. This methodology not only encourages divergent thinking in LLMs but also offers a scalable, model-agnostic approach to overcoming reasoning limitations in existing prompting strategies. Compositional Translation (CompTra) [77] is a prompting strategy designed to enhance the MT capabilities of LLMs, particularly for low-resource languages. Instead of translating entire sentences directly, CompTra treats translation as a step-by-step reasoning task. First, the input sentence is decomposed into simpler, context-preserving phrases using a prompt that includes examples from datasets like MinWikiSplit. This decomposition is dynamic and structure-aware, producing shorter phrases that are easier for LLMs to handle. Next, each phrase is translated independently in a few-shot setting, where relevant examples are retrieved through similarity search methods like BM25. Incorrectly translated phrases in other languages are filtered using a language identifier to maintain quality. Finally, the LLM is prompted again with the original sentence and the self-generated phrase-translation pairs to produce a complete, coherent translation. This approach takes advantage of the fact that LLMs perform better on shorter inputs and can more easily match relevant in-context examples for them. CompTra consistently outperforms traditional few-shot and zero-shot methods across multiple benchmarks, showing strong improvements in low-resource and out-of-domain scenarios.
4.3 Fine-Tuning Approaches
Fine-tuning involves modifying the parameters of an LLM to improve its translation capabilities. This method is computationally expensive but results in a model that is better adapted to specific translation tasks. Common fine-tuning techniques include:
4.3.1 Supervised Fine-Tuning (SFT)
Supervised fine-tuning (SFT) involves training an LLM on a large parallel corpus by minimizing the cross-entropy loss between predicted translations and ground truth references. This method significantly enhances the model’s translation accuracy, especially for specialized domains. It has been shown that SFT enables LLMs to achieve strong multilingual translation performance with as few as 32 examples, where training on one language direction generalizes to others [78]. However, using English as the target language can lead to misinterpretations, and the presence of target-side noise tends to affect well-represented languages more adversely. Ensuring successful SFT alignment requires the model to maintain a surface-level focus, preventing it from acquiring unwanted biases during the fine-tuning process. Similarly, fine-tuning improved the translation quality, and generally, large models are known to have better quality output [79]. However, while fine-tuning on parallel data significantly improves overall translation quality as measured by metrics such as COMET, it simultaneously leads to a degradation of several valuable LLM capabilities. Specifically, formality steering, few-shot domain adaptation, and document-level contextualization are all weakened as a result of continued fine-tuning. The authors observe that even with as few as 18,000 parallel training examples, these abilities begin to deteriorate, and the effect intensifies with larger datasets. These findings suggests that fine-tuning must be approached with caution.
4.3.2 Low-Rank Adaptation (LoRA)
LoRA [80] is a parameter-efficient fine-tuning technique that reduces memory overhead by injecting small trainable matrices into frozen pre-trained models. This approach enables effective adaptation to new translation domains without modifying the entire model. A comprehensive empirical study on parameter-efficient fine-tuning (PEFT) methods for MT, conducted across various parameter budgets and language pairs, observed that using adapters achieves performance on par with full fine-tuning while training only 10% of the parameters [81]. However, the effectiveness of PEFT methods varies significantly based on language pair distance and dataset size, with noticeable struggles for distant language pairs. We refer the readers to [82] and [83] for a detailed analysis for PEFT methods.
The QLoRA method [84] has been applied to fine-tune Mistral-7B [85], achieving better results than GPT-3.5-turbo when combined with one-shot prompting [86]. In a comparison between LoRA and supervised fine-tuning (SFT), both methods were found to deliver similar performance, although LoRA requires training 50 times fewer parameters [54]. Furthermore, it was demonstrated that LoRA can surpass the few-shot performance of pretrained models with as few as 2,000 training examples.
An evaluation of 15 publicly available language models on MT tasks showed that QLoRA fine-tuning outperforms both few-shot learning and models trained from scratch, achieving a significant BLEU score improvement while training on only 0.77% of the model parameters [87]. Adapter-based fine-tuning with LoRA was further shown to match the performance of traditional fine-tuning, while reducing the number of training parameters by a factor of 50 [54]. One notable application of PEFT is Language-Specific Fine-Tuning with LoRA (LSFTL) [88], which aims to improve translation quality for low-resource language pairs while maintaining computational efficiency. LSFTL selectively inserts LoRA adapters into the multi-head attention and feed-forward layers of a multilingual Transformer-based MT model. By tuning only a small number of additional parameters it avoids the high cost of full model fine-tuning. Remarkably, the approach allows models like NLLB-200-Distilled-600M to match the performance of much larger models such as NLLB-200-1.3B without LSFTL. Moreover, LSFTL shows faster convergence and better early-stage performance than full-volume fine-tuning.
4.3.3 Monolingual and Parallel data fine-tuning
A two-stage fine-tuning approach has been introduced to enhance MT capabilities in LLMs [89]. Since many LLMs are predominantly trained on English data, the first stage, known as ALMA, fine-tunes the model on non-English monolingual datasets. This step strengthens the model’s proficiency in various languages, enabling a more robust understanding of linguistic nuances across different tongues.
In the subsequent stage, ALMA fine-tunes the model using a relatively small set of high-quality parallel data. This phase focuses on translation generation, aligning the model’s outputs with accurate translations and improving its performance on translation tasks. The two-stage fine-tuning strategy has demonstrated significant improvements, achieving an average increase of over 12 BLEU and 12 COMET scores compared to the model’s zero-shot performance across 10 translation directions from the WMT’21 and WMT’22 test datasets.
To address the degradation of LLM capabilities during fine-tuning (Section 4.3.1), the authors in [79] propose a mixed-data fine-tuning strategy that incorporates both monolingual and parallel data. Their experiments demonstrate that this approach not only maintains emergent LLM capabilities but also enhances overall translation quality. For example, formality steering accuracy is preserved more effectively when monolingual data is included, and document-level contextualization shows less decline or even slight improvements compared to parallel-only fine-tuning. This finding suggests that preserving LLM strengths is not inherently at odds with achieving high translation quality. Instead, it points to the need for carefully balanced training objectives that reflect both general translation accuracy and task-specific behaviors. The use of monolingual data—especially data that structurally resembles LLM pretraining corpora—serves as an effective regularizer during adaptation, helping to retain pre-existing generalization capabilities.
4.3.4 Reinforcement Learning based Finetuning
The standard SFT methodology, which focuses on reducing the gap between predicted outputs and gold-standard references, inherently limits model performance to the quality of the training data. This presents a notable challenge, as even human-authored data—typically regarded as high-quality—can contain imperfections. To address this limitation, the authors in [90] introduced Contrastive Preference Optimization (CPO) for MT. CPO differentiates itself by training models to distinguish between “adequate” and “near-perfect” translations. This is achieved through a contrastive learning strategy that presents the model with hard negative examples—translations that are high-quality but not flawless. By learning to reject these near-perfect yet imperfect translations, the model is encouraged to produce superior outputs. When applied to ALMA models with only 22K parallel sentences and fine-tuning only 0.1% of the parameters, CPO yielded significant improvements. The authors in [91] uses Reinforcement Learning with Human Feedback (RLHF) to improve MT by aligning model outputs with human preferences. Instead of relying on expensive human-annotated preference data, the authors use high-quality human translations from books as a proxy for preferred outputs. The process involves fine-tuning a base translation model on parallel corpora, training a reward model to distinguish between machine and human translations, and then refining the translation model through RL using Proximal Policy Optimization (PPO) [92]. This approach enhances translation quality and shows cross-lingual transferability, meaning the improvements generalize to languages not directly optimized with RLHF. The results demonstrate that RLHF can effectively capture human translation preferences in a cost-efficient manner. Alternative RL-based LLM fine-tuning for MT is to used a fine-grained token-level reward [93]. xCOMET [94] identifies error spans and classifies their severity as minor, major, or critical. Then, each token in an error span is assigned a reward according to a custom severity mapping. To integrate these signals, the authors adapt standard RL algorithms into token-level variants, enabling the model to update policies at each token step rather than after full sequence generation. These token-level RL objectives provide continuous, granular feedback, improving stability and mitigating overfitting and exposure bias. The method shows significant gains across metrics when compared to sentence-level RL and supervised fine-tuning baselines. The authors in [95] propose a rule-metric mixed reward system named MT-R1-Zero that integrates both formatting constraints and automatic quality metrics to guide training. This reward combines a format reward, which enforces structured outputs using think and translate tags, with a metric reward based on established translation quality metrics — BLEU for lexical accuracy, COMETKiwi for semantic fidelity, and their combination for balanced optimization. The training leverages the GRPO algorithm, which computes advantages from multiple candidate translations to stabilize learning. MT-R1-Zero produces high-quality translation outputs, achieving strong performance across in-domain, out-of-distribution, and multilingual settings. Notably, ablation studies confirm that most of the performance gains stem from the RL framework itself, even in the absence of explicit reasoning prompts, highlighting the efficacy of the rule-guided, reward-driven optimization process.
5 LLMs and Encoder-Decoder across Languages
This section explores how LLMs perform across languages and how they compare to dedicated MT systems. An early study found that ChatGPT lags behind traditional MT systems such as Google Translate, DeepL, and Tencent, but observed that GPT-4 performs competitively with these systems in zero-shot settings [48]. This indicates that some LLMs achieve performance on par or even better with traditional MT systems for high-resource languages.
When translating from Spanish to 11 indigenous South American languages, fine-tuned M2M100 was shown to consistently outperform GPT-4 [96]. The performance gap further increases when M2M100 is combined with techniques like kNN-MT [97]. The contrast between high-resource and low-resource languages is also highlighted by a multitask evaluation, where ChatGPT’s performance, although slightly lower, remains close to fine-tuned state-of-the-art models [98]. For instance, ChatGPT achieves 58.64 chrF++ compared to 63.5 for fine-tuned models in XXEn translation, and 51.12 chrF++ versus 54.4 for EnXX [44]. However, for low-resource languages, the difference becomes much more pronounced, with ChatGPT scoring 27.79 chrF++ compared to 54.9 for fine-tuned models in XXEn, and 21.57 versus 41.9 chrF++ for EnXX [69]. A similar study [99] identifies key bottlenecks, such as the asymmetry in translation quality between high-resource to low-resource directions and the poor generative capability of smaller models. The study finds that LLMs generally underperform for low-resource languages compared to traditional encoder-decoder NMT systems, which is also confirmed by [100] earlier. Further, they observe that generating En XX is more difficult compare to generating XXEn.
Given the poor state of performance in low-resource languages, the authors in [101] examine whether LLMs can learn to translate extremely low-resource languages using only grammar books. They split grammar books into parallel examples and grammatical explanations, then tested LLMs and a fine-tuned NMT model on translation, grammatical judgment, and gloss prediction tasks. They also introduced typological prompts summarizing high-level grammatical features. Their findings show that translation improvements come almost entirely from parallel examples, not grammatical explanations. Fine-tuned models performed as well or better than LLMs using grammar books. However, for linguistic tasks like grammaticality judgment and gloss prediction, typological prompts combined with parallel examples led to strong performance. The study concludes that for translation, collecting parallel data is more effective than relying on descriptive grammar, while structured grammatical knowledge is better suited for non-translation tasks. The authors in [102] investigates the persistent issue of the Curse of Multilinguality (CoM) in LLM-based multilingual MT, where scaling to more languages introduces performance bottlenecks. The authors identify a key phenomenon: asymmetry in linguistic conflicts and synergies—with translation from other languages to English (XXEn) suffering from conflicts, while English to other languages (EnXX) benefits from synergy. These issues primarily arise during the post-training phase rather than during multilingual pretraining. To address this, they propose a Direction-Aware Training (DAT) approach: applying separate training for XXEn directions to reduce conflicts and group multilingual training for EnXX to leverage synergy. They further introduce group-wise model merging (DATM) to improve scalability by reducing the number of LoRA adapters while maintaining translation quality. This work highlights the importance of targeted post-training strategies and provides a path toward building high-quality, cost-effective MT systems.
Towards building efficient LLM-based MT systems, a pre-trained decoder-only LLM can serve as the encoder, paired with a lightweight NMT-style decoder [103]. To enable compatibility between the LLM’s output and the decoder’s input, the intermediate representations from multiple LLM layers are integrated, transformed to a lower dimensional space, and optionally enhanced with additional encoder layers to capture bidirectional context. Training proceeds in two phases: initially, only the transformation layers and decoder are trained while keeping the LLM frozen; subsequently, the entire model is fine-tuned on ComMT, a comprehensive benchmark covering general, domain-specific, document-level, terminology-constrained, and post-editing translation tasks. This approach, called LaMaTE, demonstrates strong performance and generalization, while significantly improving efficiency—achieving up to 6.5× faster decoding and 75% reduction in memory usage—highlighting the potential of LLMs as encoders in NMT architectures.
A key distinction between LLMs and encoder-decoder models is their ability to generalize beyond sentence-level translation, offering rich contextualization, stylistic control, and flexibility via prompting. A recent study [79] reveals that this distinction erodes when LLMs are fine-tuned on parallel data. In many respects, fine-tuned LLMs begin to resemble traditional NMT systems: they lose control over output formality, struggle with context-sensitive disambiguation, and fail to generalize effectively to unseen domains without explicit retraining. While idiomatic translation often improves—particularly when initial performance is weak—other capabilities, such as formality control and document-level contextualization, suffer significant degradation. This degradation is not readily apparent when relying solely on standard evaluation metrics. For instance, COMET scores may remain stable or even improve while the model’s ability to follow stylistic constraints or disambiguate based on discourse context declines. This indicates a mismatch between what current MT metrics measure and what users value in real-world translation tasks. This convergence suggests that careless adaptation of LLMs risks negating their primary benefits. While LLMs initially surpass NMT systems in their versatility, indiscriminate fine-tuning can narrow their capabilities to those of more brittle translation architectures. This finding reinforces the need to develop fine-tuning strategies that preserve the emergent capabilities of LLMs rather than inadvertently discarding them.
6 LLM based Evaluation
The performance of LLMs) on the MQM-2020 dataset has been evaluated by comparing their Spearman correlations with traditional MT metrics, such as ROUGE, BERTScore, and BARTScore [104]. The findings indicate that, when provided with appropriate instructions and demonstrations, LLMs can outperform these conventional metrics. However, the comparison is restricted to high-resource language pairs (English and Chinese), and more advanced metrics, such as COMET, which exhibit higher correlation with human judgments, were not included.
For MT evaluation, GEMBA-style prompting [105] has been employed, showing that GPT-4 achieves the highest system-level pairwise accuracy and the highest Kendall’s tau [106] for English-to-German (EnDe) and English-to-Russian (EnRu) translations [107]. In contrast, COMET demonstrates superior performance for Chinese-to-English (ZhEn), achieving the highest Kendall’s tau correlation.
Building on these approaches, Error Analysis Prompting has been introduced to combine error analysis with CoT reasoning for MT evaluation using LLMs [108]. Experiments reveal that, while LLMs achieve higher system-level accuracy, they fall short of traditional metrics, including embedding-based approaches such as BLEURT, COMET, and MetricX, at the segment level.
The authors in [111] investigates the effectiveness of LLM in evaluating MT quality without requiring fine-tuning. The authors systematically examine what translation-related information, such as the source sentence, reference translation, identified errors, and annotation guidelines is necessary for LLMs to reliably assess MT quality. The findings highlight that the inclusion of both the source and reference translations is crucial for improving evaluation accuracy, while error annotations and detailed guidelines offer limited additional benefit. Larger models tend to benefit more from CoT prompting, though smaller models also perform competitively. However, LLMs are inconsistent in generating numerical scores, raising concerns about their reliability for automated MT evaluation. For a more robust evaluation, M-MAD [112] use a three-stage framework for via structured multi-agent collaboration. First, it decouples the MQM evaluation guideline into four distinct dimensions - accuracy, fluency, style, and terminology — allowing dimension-specific LLM agents to identify and annotate errors independently. In the second stage, each dimension undergoes a focused two-agent debate, where one agent supports and the other critiques the initial annotation across multiple rounds, fostering more accurate and balanced error assessments. Finally, a judge agent synthesizes the debated viewpoints and computes the overall MQM score using weighted counts of major and minor errors. This structured approach significantly enhances system-level correlation; however, non-LLM metrics fared better in segment-level.
7 Summary and Recommendation
This survey has provided a comprehensive analysis of how LLMs are reshaping the landscape of MT. It highlighted the effectiveness of LLMs in generating synthetic data, leveraging few-shot and in-context learning, and enhancing performance through parameter-efficient fine-tuning techniques such as LoRA and QLoRA. While LLMs demonstrate competitive results in high-resource language pairs, they continue to underperform compared to traditional MT systems in low-resource settings due to persistent challenges like data scarcity and linguistic diversity. Furthermore, the evaluation of LLM-based MT reveals that despite achieving strong system-level accuracy, LLMs often struggle with segment-level consistency, as measured by metrics such as COMET. Despite the progress, a number of open challenges persist, motivating future research directions as outlined below:
Enhanced Adaptation for Low-Resource Languages: Despite advancements, LLMs struggle with low-resource languages. Future research should focus on improved transfer learning strategies, cross-lingual pretraining, and multilingual fine-tuning to enhance performance in low-data settings. Leveraging smaller, domain-specific datasets and semi-supervised learning could boost generalization in low-resource scenarios.
Reinforcement Learning for Quality Control: The application of reinforcement learning techniques (e.g., RLHF or Contrastive Preference Optimization) can refine LLMs’ output by aligning them with human preferences. This would reduce hallucination issues and improve translation reliability, particularly in domain-specific and low-resource contexts.
Better Evaluation and Interpretability: Future work should explore more robust and interpretable evaluation metrics. Incorporating explainability techniques, such as attention visualization or error analysis prompting, could help in understanding LLM translation errors and guiding model improvements.
It is important to emphasize that MT with LLMs is a vast and rapidly evolving area of research. Each aspect—be it data generation, learning paradigms, fine-tuning strategies, or evaluation—warrants a dedicated survey of its own. Due to space and scope limitations, this review focuses primarily on sentence-level translation. Consequently, while we aimed to cover a representative subset of the literature, some relevant works may not have been included.
Acknowledgment
Baban Gain and Dibyanayan Bandyopadhyay gratefully acknowledge the Prime Minister’s Research Fellowship (PMRF) program for providing financial support and enabling this research.